# An analytical exploration of the optomechanical attractor diagram and of limit cycles

### v1

@@ Abstract
We analyse the interplay between mechanical and radiation pressure in an optomechanical cavity system. Our study is based on an analytical evaluation of the infinite Bessel summations involved, which previously had led to a numerical exploration of the so-called attractor diagram. The analytical expressions are then suitable for further asymptotic analysis in opposing regimes of the amplitude, which allows for a characterisation of the diagram in terms of elementary functions. The same method is applied to the study of optomechanical limit cycles, extending previous results for enhanced values of the parameter space without the need for approximations such as the resolved sideband approximation.
@@ I Introduction
In the initial section, we present a methodology for the analytical study of the ensuing attractor diagram that has been numerically outlined in references [7, 8]. The determination of radiation pressure represents a fundamental problem in optomechanical systems and the force and power input are well-known to be characterised by infinite sums of Bessel functions [7, 3]. As we will see at the end of the Outlook section, these infinite summations are ubiquitous in a wide range of problems in cavity optomechanics and magnomechanics and beyond, including an extremely large number of quantum Floquet problems. This paper will exploit that these sums can be actually computed exactly [9, 10] and discuss a number of interesting predictions from the exact formulas, focusing exclusively on the dynamical multistability [7, 8, 11] and limit cycle problems in cavity optomechanics [12, 13].
@@
One of the applications of the effective equation of motion is the subsequent description of optomechanical limit cycles in the quantum regime. The objective of this paper as it concerns to limit cycles is to characterise the drift and diffusion coefficients in [12] again through the utilisation of exact evaluations of Bessel summations, previously employed in [9, 10]. The results will be compared with those presented in [12], which did not consider the exact summation of the formulas. This led them to consider a drive setting close to resonance, specifically in the sideband-resolved regime with a detuning close to the resonance. This resonance implies that only the first terms in the sums need to be retained. We investigate the differences and further possibilities afforded by considering the summation in full, analytically. For example, some limit cycles that appear in [12] turn out to be an artifact of the approximation as they disappear as solutions in the exact formula. We will also specifically examine the implications of the exact summation of the diffusion coefficient for the Wigner function.
@@ II Radiation pressure: parametric instability and Cooling
In the last decades the classical theory of the optomechanical parametric instability produced by radiation pressure backaction has been extensively studied in a large number of works, starting with the influential experimental works [17, 4]. We shall follow the very well-known ensuing theoretical analysis of [7, 11], where a numerical analysis of the analytical expressions of the model were carried out, culminating in a rather pervasive attractor diagram, that has also been portrayed in many reviews and works afterwards [18]. A consequence of our analytical formulas obtained below will be a quantitative analysis of this attractor diagram for certain regions of the parameter space. These results apply to rather generic optomechanical systems, consisting of a driven optical cavity and a movable mirror attached to a cantilever.
The analysis in [7] relied on a simple ansatz to determine the cavity’s amplitude $A$. In particular,
@@
for the self-induced oscillations. The radiation pressure force $F(t)$ is deduced by solving the classical equation for the light field amplitude, which is
@@
The time-averaged power input due to this force is $\left\langle F\dot{x}\right\rangle$. After inserting the ansatz (1), the solution for the optical degrees of freedom can be obtained as a Fourier series $\alpha(t)=e^{i\varphi(t)}\sum_{n}\alpha_{n}e^{in\Omega_{{\rm m}}t}$, with coefficients [7]
@@ II.1 Analytical characterization
These series have always been dealt with numerically. Following the same analytical results employed while studying superconducting qubits [9] (see the Appendix for a summary), we can now give their full analytical form. For the time-averaged force we obtain
@@
Similarly as in the previous examples, one can give the asymptotic formulas at $|X|\gg 1$, $|X|\gg|\nu|^{2}$. In particular, we find for the force average:
@@
Thus the functional dependence of the time-averaged force satisfies properties for special values of the parameters. In particular, it is worth noticing the change of periodicity in $r$ occurring at $2GA=n\pi\Omega_{{\rm m}}$, $n\in\mathbb{Z}$, the suppression of oscillations in $A$ occurring at $2GA=(n+\frac{1}{2})\pi\Omega_{{\rm m}}$ and the expected resonances occurring at integer values of $r$ and small $c$.
@@ II.2 Parametric instability, amplification and cooling: The attractor diagram
The condition of force balance determines the average position of the oscillator, thereby yielding an implicit equation for $\bar{x}$ [7],
$\langle\ddot{x}\rangle\equiv 0\quad\Leftrightarrow\quad m\omega_{M}^{2}\bar{x}=\left\langle F_{{\rm rad}}\right\rangle=\frac{\hbar g}{mx_{\text{ZPF}}}\langle\left|\alpha(t)\right|^{2}\rangle\,,$ (14)
where the average radiation force, $\langle F_{\text{rad}}\rangle$ is a function of the parameters $\bar{x}$ and $A$.
@@
For each value of the oscillation amplitude $A$ the plot of the ratio between radiation power input and friction loss, $P_{{\rm rad}}/P_{{\rm fric}}=\langle F_{\text{rad}}\dot{x}\rangle/(\Gamma_{M}\langle\dot{x}^{2}\rangle)$ is analyzed in [7, 11]. Therefore, the study of the ratio has been subjected to graphical and numerical analysis as has been also reviewed on numerous occasions after [7, 11]. The power balance is fulfilled if the ratio is one, whereas if the power input into the cantilever by radiation pressure is greater than frictional losses (i.e., for a ratio greater than one), the amplitude of oscillations will increase, otherwise it will decrease. Thus, stable solutions (dynamical attractors) are given by decrements in the ratio while increasing oscillation amplitude (energy).
@@ II.2.1 Large driving amplitudes and resolved/unresolved sideband limit
which shows that the dependence with the amplitude $X$ is of zero order, whereas for the friction pressure is always quadratic. Therefore, for very large amplitudes mechanical friction pressure will always dominate and the ratio goes to 00, as the attractor diagram also clearly shows. Intermediate amplitudes are still well described by the asymptotic formula (16) if $\mu$ is not too large. Notice also that, in general, for very large amplitudes, the starting Ansatz itself (1) is not correct anyway [7].
In [7, 11], it is observed that for sufficiently strong driving, self-induced oscillations appear around integer multiples of the cantilever frequency, $\Delta\approx n\omega_{M}$. Notice that our solution is generically oscillatory both in the detuning $r$ and the amplitude $X$ (the only dependence on the amplitude) for all of its domain of validity. The denominator term $(\cos(2\pi r)-\cosh(2\pi c))$ is always negative hence the overall sign depends on the $\sin(\pi r)$ and $\cos(2X)$ terms.
@@
The attractor diagram in [7, 11] has been studied for the resolved sideband regime. This regime is of interest because it is crucial for ground-state cooling [8, 21], and $c=1/20$ has already been realised experimentally in [22]. For $r=(2m+1)/4$ and $m=0,1,2,...$ the two sideband resolution limits give (for simplicity we write the case $m=0$):
$\left\langle F\dot{x}\right\rangle$ $\simeq$ $\frac{\kappa^{2}\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{4G}\frac{\sqrt{2}e^{-\pi c}\cos(2X)}{c}\text{ for }c\gg 1,$
$\left\langle F\dot{x}\right\rangle$ $\simeq$ $\frac{\kappa^{2}\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{4G}4\sqrt{2}c\cos(2X)\text{ for }c\ll 1.$
The functional dependence with $c$ and $X$ is the same for other choices of $r$ that do not vanish the pressure (that is, other than $r$ integer). For example, for $r=1/2$, then:
$\left\langle F\dot{x}\right\rangle$ $\simeq$ $\frac{\kappa^{2}\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{4G}\frac{2e^{-\pi c}\cos(2X)}{c}\text{ for }c\gg 1,$
$\left\langle F\dot{x}\right\rangle$ $\simeq$ $\frac{\kappa^{2}\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{4G}4c\cos(2X)\text{ for }c\ll 1.$
@@ II.2.2 Small driving amplitudes
It follows immediately by having the Bessel function evaluation. Using [Watson, 5.41, Eq1] we have:
$J_{\nu}(x)J_{\mu}(x)=\sum_{m=0}^{\infty}\frac{(-1)^{m}\left(\frac{x}{2}\right)^{\mu+\nu+2m}\left(\mu+\nu+m+1\right)_{m}}{m!\Gamma\left(\nu+m+1\right)\Gamma\left(\mu+m+1\right)}.$
At first order, it is linear in $x$ and that makes the radiation pressure quadratic in the amplitude, for small amplitudes as expected. We have at first order:
@@
Cooling occurs when the ratio is negative since then the radiation pressure is negative. Recall that $c>0$ then the negative-pressure condition is simply
@@
$\frac{1}{\Omega_{\mathrm{m}}}\left(G\overline{x}+\Delta\right)<0\ .$
Let us now consider the implications of the bound (17). A close examination of the function $f(r)$ shows that the maximum value occurs for $c=1/\sqrt{2}$ at $r=c$, where
$f(x_{\rm max})\bigg{|}_{c_{\rm max}}=\frac{G\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{4\pi\Omega_{\mathrm{m}}}<1$ (18)
@@
Figure 1: The maximum value of the ratio $f(r)\equiv\frac{\left\langle F\dot{x}\right\rangle}{P_{fric}}$ as a function of $c$ (in units of $2G\big{|}\alpha_{\mathrm{max}}\big{|}^{2}/\pi\Omega_{\mathrm{m}}$).
@@
$f(r_{0})\bigg{|}_{c\ll 1}=\frac{G\big{|}\alpha_{\mathrm{max}}\big{|}^{2}}{2\pi\Omega_{\mathrm{m}}}\ c<1$ (19)
@@
The equation obtained implies that for blue-detuning $\Delta>0$ one could also have cooling is the mechanical off-set $\overline{x}$ is such that $G\overline{x}<-\Omega_{\mathrm{m}}\Delta$. If we discuss, as in [11], in terms purely of the detuning $\Delta$ with the off-set $\overline{x}=0$. Then the condition is simply that red-detuning  [^1] For red-detuning $\Delta<0$ one can typically have cooling. To be more precise, the theory of resonator dynamics indicates that in this regime the energy tends to be transferred from the resonator to the cavity [7]. The back-action damping is positive, guaranteeing that the system will remain in a state where the mechanical oscillator fluctuates about a fixed point. leads to cooling whereas blue-detuning to amplification, as usual.

### v2

# Analytical exploration of the optomechanical attractor diagram and of limit cycles

@@ Abstract
We analyse the interplay between mechanical and radiation pressure in an optomechanical cavity system. Our study is based on an analytical evaluation of the infinite Bessel summations involved, which previously had led to a numerical exploration of the so-called attractor diagram. The analytical expressions are then suitable for further asymptotic analysis in opposing regimes of the amplitude, which allows for a characterisation of the diagram in terms of elementary functions. Building on this framework, we investigate the emergence and properties of optomechanical limit cycles beyond the constraints of the resolved sideband approximation. By employing a Fokker-Planck formalism originally developed in the context of laser theory and then used in cavity optomechanics, we describe the quantum regime of these limit cycles, offering a more detailed and unified analytical perspective.
@@ I Introduction
More precisely, we begin with the work carried out on a standard optomechanical system [7, 8], which involves a high-finesse optical cavity formed by two mirrors: one fixed and the other mounted on a flexible cantilever. The position of the movable mirror, $x$, affects the cavity length, thereby shifting the resonance frequency and modulating the amplitude of the light inside the cavity, $\alpha$. The radiation pressure force, proportional to $|\alpha|^{2}$, acts back on the mirror to create a coupled optomechanical system. This nonlinear coupling gives rise to complex dynamical phenomena, including multistability and self-sustained oscillations. For certain parameters, the static force balance has multiple stable points, leading to bistability. The combination of radiation pressure delays and mechanical damping creates dynamical instabilities, causing self-sustained oscillations of the cantilever. The system exhibits a range of stable oscillation amplitudes (dynamic attractors) that depend on the detuning, $x_{0}$ and the ratio between the mechanical and radiation pressure $\Gamma/P$.
In the next section, we present a methodology for the analytical study of the expression for the radiation pressure in the above described model and then the ensuing attractor diagram that has been numerically outlined in references [7, 8]. The determination of radiation pressure represents a fundamental problem in optomechanical systems and the force and power input are well-known to be characterised by infinite sums of Bessel functions [7, 3], as shown in the next Section. Likewise, as we will see at the end of the Outlook section, these infinite summations are ubiquitous in a wide range of problems in cavity optomechanics and magnomechanics and beyond, including an extremely large number of quantum Floquet problems. This paper will exploit that these sums can be actually computed exactly [9, 10] and discuss a number of interesting predictions from the exact formulas, focusing exclusively on the dynamical multistability [7, 8, 11] and limit cycle problems in cavity optomechanics [12, 13].
@@
One of the applications of the effective equation of motion is the subsequent description of optomechanical limit cycles in the quantum regime. The objective of this paper as it concerns to limit cycles is to characterise the drift and diffusion coefficients in [12] again through the utilisation of exact evaluations of Bessel summations, previously employed in [9, 10]. The results will be compared with those presented in [12], which did not consider the exact summation of the formulas. This led the authors of that work to consider a drive setting close to resonance, specifically in the sideband-resolved regime with a detuning close to the resonance. This resonance implies that only the first terms in the sums need to be retained. We investigate the differences and further possibilities afforded by considering the summation in full, analytically. For example, some limit cycles that appear in [12] turn out to be an artifact of the approximation as they disappear as solutions in the exact formula. We will also specifically examine the implications of the exact summation of the diffusion coefficient for the Wigner function.
@@ II Radiation pressure: parametric instability and Cooling
In the last decades the classical theory of the optomechanical parametric instability produced by radiation pressure backaction has been extensively studied in a large number of works, starting with the influential experimental works [17, 4]. We shall follow the very well-known ensuing theoretical analysis of [7, 11], where a numerical analysis of the analytical expressions of the model were carried out, culminating in a rather pervasive attractor diagram, that has also been portrayed in many reviews and works afterwards [18].
The model studied in these works describe the nonlinear dynamics of an optical cavity in which one mirror is mounted on a flexible mechanical element (a cantilever). The cavity is driven by a laser, and the radiation pressure exerted by the cavity light couples the mechanical motion of the mirror to the intracavity optical field. A consequence of our analytical formulas obtained below will be a quantitative analysis of this attractor diagram for certain regions of the parameter space. These results apply to rather generic optomechanical systems, consisting of a driven optical cavity and a movable mirror attached to an oscillator, see the fig. 1.
Figure 1: Generic optomechanical setting with detuning $\Delta=\omega_{L}-\omega_{c}$ and $G=\partial\omega_{c}/\partial x$ (frequency shift of the cavity resonance per unit mechanical displacement). The classical coupled equations of motion are given below in (1) and (2).
Writing $\alpha(t)$ for the slowly varying complex amplitude of the intracavity optical field, which is the classical equivalent of the cavity photon field operator $\alpha(t)=\left\langle\widehat{a}(t)\right\rangle$, the coupled classical nonlinear dynamics of the system are governed by the following coupled equations of motion for $\alpha(t)$ and for the mechanical displacement of the oscillator from its equilibrium position $x(t)=\left\langle\widehat{x}(t)\right\rangle$ [7, 11, 3]:
$\dot{\alpha}=-\frac{\kappa}{2}\alpha+i(\Delta+Gx)\alpha+\sqrt{\kappa_{ex}}\alpha_{in},$ (1)
$m\ddot{x}=-m\Omega_{m}^{2}x-m\Gamma_{m}\dot{x}+\hbar G|\alpha|^{2}+F_{ext},$ (2)
where $\kappa$ denotes the total energy decay rate of the optical cavity mode and describes the rate at which energy leaks out of the optical cavity. The parameter $\Delta$ is the detuning of the driving laser frequency $\omega_{L}$ from the bare cavity resonance frequency $\omega_{c}$, defined as $\Delta=\omega_{L}-\omega_{c}$. $G$ is the optomechanical coupling strength describing the frequency shift of the cavity resonance per unit mechanical displacement and $\alpha_{in}$ is the input field amplitude (drive) of the laser incident on the cavity. The effective mass of the mechanical oscillator is $m$ and the intrinsic mechanical resonance angular frequency is $\Omega_{m}$. The mechanical energy decay rate (damping rate), describing mechanical losses is denoted by $\Gamma_{m}$. Finally, $|\alpha|^{2}$ denotes the intracavity photon number, i.e. the intensity of the optical field inside the cavity and $F_{ext}(t)$ denotes any other external force acting on the mechanical oscillator (e.g. thermal noise or applied forces). We adopted the notation of the review [3].
We have adopted the notation in [3], also for the Ansatz below and the ensuing solution, described below. Notice that in the original works [7, 11] the parameters are presented in a seemingly different yet equivalent manner.
These coupled equations capture the dynamics of cavity optomechanics in the classical limit, describing how light inside the cavity influences the motion of the mechanical oscillator via radiation pressure and how the oscillator, in turn, shifts the cavity resonance frequency and affects the optical field. More precisely, in (1) the cavity field amplitude $\alpha$ decays at a rate $\kappa/2$, acquires a phase shift depending on the detuning $\Delta$ and mechanical displacement via $Gx$, and is driven by the external input field $\alpha_{in}$ through the coupling rate $\kappa_{ex}$ and in (2) the mechanical oscillator behaves like a damped harmonic oscillator with mass $m$, resonance frequency $\Omega_{m}$, damping $\Gamma_{m}$, and is driven by the radiation pressure force $\hbar G|\alpha|^{2}$ (force from the intracavity photons) plus any other external force $F_{ext}$.
The Ansatz for the self-induced oscillations is [7]
@@
The radiation pressure force $F(t)$ is deduced by solving the classical equation for the light field amplitude, which is
@@
The time-averaged power input due to this force is $\left\langle F\dot{x}\right\rangle$. After inserting the ansatz (3), the solution for the optical degrees of freedom can be obtained as a Fourier series $\alpha(t)=e^{i\varphi(t)}\sum_{n}\alpha_{n}e^{in\Omega_{{\rm m}}t}$, with coefficients [3, 7]
@@ II.1 Analytical characterization
These series have always been dealt with numerically. Following the same analytical results employed while studying superconducting qubits [9] (see Appendix B for a summary), we can now give their full analytical form. For the time-averaged force we obtain
@@
Similarly as in the previous examples, one can give the asymptotic formulas at $|X|\gg 1$, $|X|\gg|\nu|^{2}$ using (37). In particular, we find for the force average:
@@
From these formulas one can see that the functional dependence of the time-averaged force satisfies properties for special values of the parameters:
  • When $r=n+\frac{1}{2}$, with integer $n$, one has $\cos(\pi r)=0$ so the coefficient of the $\sin(\frac{2GA}{\Omega_{{\rm m}}})$ vanishes in $\langle F\rangle$. Therefore oscillations of $\langle F\rangle$ as a function of $A$ are strongly suppressed at large $A$ ($\langle F\rangle$ then has smaller oscillations of $O(1/A^{2})$ originating from subleading corrections to the asymptotic formula (37) of the Bessel function).
  • $\langle F\rangle$ is a periodic function of $r$ with periodicity $r\to r+2$. However, at the special values $2X=2GA/\Omega_{n}=\pi n$, with integer $n$, one has $\sin(\frac{2GA}{\Omega_{{\rm m}}})=0$ and the periodicity of $\langle F\rangle$ changes to $r\to r+1$.
  • For small $\kappa$, and $r$ generic, one has $\langle F\rangle\propto\kappa^{2}$. However, at integer values of $r$, $\langle F\rangle$ exhibits resonance behavior, with peaks values given by
    $\langle F\rangle\bigg|_{r=n}=\frac{|\alpha_{\rm max}|^{2}\Omega_{m}}{\pi A}\left(1-(-1)^{n}\sin(\frac{2GA}{\Omega_{{\rm m}}})\right)\ .$ (16)
    Thus, for integer values of $r$, the small factor $\kappa^{2}$ cancels out against a factor $1/\kappa^{2}$ originating from the resonance. A similar resonance behavior is seen in the exact formula (12) for $\langle F\dot{x}\rangle$. For generic $r$ is $O(\kappa^{3})$ but at integer $r$ becomes $O(\kappa)$.
The first two properties arise exclusively as a result of the exact resummation carried out in the present study. It should be feasible to experimentally test these remarkable predicted features.
@@ II.2 Parametric instability, amplification and cooling: The attractor diagram
The condition of force balance determines the average position of the oscillator, thereby yielding an implicit equation for $\bar{x}$ [7, 11, 3],
$\langle\ddot{x}\rangle\equiv 0\quad\Leftrightarrow\quad m\omega_{M}^{2}\bar{x}=\left\langle F_{{\rm rad}}\right\rangle=\hbar G\langle\left|\alpha(t)\right|^{2}\rangle.$ (17)
The average radiation force, $\langle F_{\text{rad}}\rangle$ is a function of the parameters $\bar{x}$ and $A$.
@@
For each value of the oscillation amplitude $A$ the plot of the ratio between radiation power input and friction loss, $P_{{\rm rad}}/P_{{\rm fric}}=\langle F_{\text{rad}}\dot{x}\rangle/(\Gamma_{M}\langle\dot{x}^{2}\rangle)$ is analyzed in [7, 11]. Therefore, the study of the ratio has been subjected to graphical and numerical analysis as has been also reviewed on numerous occasions after [7, 11]. The power balance is fulfilled if the ratio is 1, whereas if the power input into the cantilever by radiation pressure is greater than frictional losses (i.e., for a ratio greater than 1), the amplitude of oscillations will increase, otherwise it will decrease. Thus, stable solutions (dynamical attractors) are given by decrements in the ratio while increasing oscillation amplitude (energy).
@@ II.2.1 Large driving amplitudes and resolved/unresolved sideband limit
which shows that the dependence with the amplitude $A$ is oscillating but with no power dependence on $A$. This may be compared with the friction pressure $P_{{\rm fric}}=\Gamma_{M}\left\langle\dot{x}^{2}\right\rangle$, which is always quadratic (see (3)). Therefore, for very large amplitudes mechanical friction pressure will always dominate and the ratio goes to 0, as the attractor diagram also clearly shows (we refer the reader to Figure 2 of [11], paying particular attention to the floor of the plot region, which holds for all values of detuning at higher energies). Intermediate amplitudes are still well described by the asymptotic formula (19) if $c$ is not too large. Notice also that, in general, for very large amplitudes, the starting Ansatz itself (3) is not correct anyway [7].
In [7, 11], it is observed that for sufficiently strong driving, self-induced oscillations appear around integer multiples of the cantilever frequency, $\Delta\approx n\omega_{M}$. Notice that our solution is generically oscillatory both in the detuning $r$ and the amplitude $X$ (the only dependence on the amplitude) for all of its domain of validity. The denominator term $(\cos(2\pi r)-\cosh(2\pi c))$ is always negative and hence the overall sign depends on the $\sin(\pi r)$ and $\cos(2X)$ terms.
@@
The attractor diagram in [7, 11] has been studied for the resolved sideband regime. This regime is of interest because it is crucial for ground-state cooling [8, 21], and $c=1/20$ has already been realised experimentally in [22]. Using (19), for $r=(2m+1)/4$ and $m=0,1,2,$... the two sideband resolution limits give (for simplicity we write the case $m=0$):
$\left\langle F\dot{x}\right\rangle$ $\simeq$ $\frac{\kappa^{2}\big|\alpha_{\mathrm{max}}\big|^{2}}{4G}\frac{\sqrt{2}e^{-\pi c}\cos(2X)}{c}\quad\text{ for }c\gg 1,$
$\left\langle F\dot{x}\right\rangle$ $\simeq$ $\frac{\kappa^{2}\big|\alpha_{\mathrm{max}}\big|^{2}}{4G}4\sqrt{2}c\cos(2X)\quad\text{ for }c\ll 1.$
The functional dependence on $c$ and $X$ remains unchanged for those values of $r$ that do not lead to vanishing pressure –that is, for non-integer values of $r$. For example, for $r=1/2$, then:
$\left\langle F\dot{x}\right\rangle$ $\simeq$ $\frac{\kappa^{2}\big|\alpha_{\mathrm{max}}\big|^{2}}{4G}\frac{2e^{-\pi c}\cos(2X)}{c}\quad\text{ for }c\gg 1,$
$\left\langle F\dot{x}\right\rangle$ $\simeq$ $\frac{\kappa^{2}\big|\alpha_{\mathrm{max}}\big|^{2}}{4G}4c\cos(2X)\quad\text{ for }c\ll 1.$
@@ II.2.2 Small driving amplitudes
It follows immediately by having the Bessel function evaluation (12). Using [26] (5.41, Eq1) we have:
$J_{\nu}(x)J_{\mu}(x)=\sum_{m=0}^{\infty}\frac{(-1)^{m}\left(\frac{x}{2}\right)^{\mu+\nu+2m}\left(\mu+\nu+m+1\right)_{m}}{m!\Gamma\left(\nu+m+1\right)\Gamma\left(\mu+m+1\right)},$
where $(n)_{m}$ denotes the rising factorial (or Pochhammer symbol). At first order, it is linear in $x$ and that makes the radiation pressure quadratic in the amplitude, for small amplitudes as expected. We have at first order:
@@
Cooling occurs when the ratio is negative since then the radiation pressure is negative. Recall that $c>0$ and then the negative-pressure condition is simply
@@
$\frac{1}{\Omega_{\mathrm{m}}}\left(G\overline{x}+\Delta\right)<0\ .$ (21)
Let us now consider the implications of the bound (20). A close examination of the function $f(r)$ shows that the maximum value occurs for $c=c_{\rm max}=1/\sqrt{2}$ at $r=r_{\rm max}=c$, where
$f(r_{\rm max})\bigg|_{c_{\rm max}}=\frac{G\big|\alpha_{\mathrm{max}}\big|^{2}}{4\pi\Omega_{\mathrm{m}}}<1\ .$ (22)
@@
Figure 2: The maximum value of the ratio $f(r)\equiv\frac{\left\langle F\dot{x}\right\rangle}{P_{fric}}$ as a function of $c$ (in units of $2G\big|\alpha_{\mathrm{max}}\big|^{2}/\pi\Omega_{\mathrm{m}}$). The dashed lines correspond to the small $c$ and large $c$ behavior, governed by the formulas (23) and (24), respectively.
@@
$f(r_{0})\bigg|_{c\ll 1}=\frac{G\big|\alpha_{\mathrm{max}}\big|^{2}}{2\pi\Omega_{\mathrm{m}}}\ c<1\ .$ (23)
@@
The equation (21) obtained implies that for blue-detuning $\Delta>0$ one could also have cooling is the mechanical off-set $\overline{x}$ is such that $G\overline{x}<-\Delta$. If we discuss, as in [11], in terms purely of the detuning $\Delta$ with the off-set $\overline{x}=0$, then the condition is simply that red-detuning  [^1] For red-detuning $\Delta<0$ one can typically have cooling. To be more precise, the theory of resonator dynamics indicates that in this regime the energy tends to be transferred from the resonator to the cavity [7]. The back-action damping is positive, guaranteeing that the system will remain in a state where the mechanical oscillator fluctuates about a fixed point. leads to cooling whereas blue-detuning to amplification, as usual.